40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 8 15:31:48 2013, Jenne, Update, ASC, POP QPD calibration prep 
    Reply  Tue Jul 9 11:41:22 2013, Jenne, Update, ASC, POP QPD calibration attempt 
       Reply  Tue Jul 9 16:08:32 2013, Jenne, Update, ASC, POP QPD calibration attempt 
          Reply  Fri Jul 12 21:23:42 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeVertMicrometer_July2013.pngChangeSideMicrometer_July2013.png
             Reply  Mon Jul 15 17:16:59 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeSideMicrometer_July2013_calib2.pngChangeSideMicrometer_July2013_calib1.pngChangeVertMicrometer_July2013_calib1.png
                Reply  Tue Jul 16 01:17:21 2013, Jenne, Update, ASC, POP QPD calibration attempt 
                   Reply  Tue Jul 23 01:30:27 2013, Jenne, Update, ASC, POP QPD analysis 9x
                      Reply  Fri Jul 26 13:39:30 2013, Koji, Update, ASC, POP QPD analysis 
Message ID: 8810     Entry time: Tue Jul 9 11:41:22 2013     In reply to: 8805     Reply to this: 8812
Author: Jenne 
Type: Update 
Category: ASC 
Subject: POP QPD calibration attempt 

I was bad, and forgot to elog the most important part of my work yesterday - that I had rotated the POP QPD by 90 degrees, so that I could fit the micrometer onto the table.  There is a sticker on the front of the QPD to indicate which direction is "X" and "Y" for the output of the readout box.  Right now (and the way that I will mount the QPD to the table, after I redo the calibration today), X is PITCH, and Y is YAW.  Koji and Nic swapped the cables to the ADC to make this all consistent.

Yesterday, I locked the PRM-ITMY half cavity, and tried to take calibration data.  However, with no ASC servo engaged, the beam was still moving.  Also, with only the half-cavity, I had very little light on the QPD, and since it has internal normalization, the outputs can get a little funny if there isn't enough light.  I had checked, and even with the gain cranked up to maximum, the "light level too low" LED was illuminated.  So, my calibration data from yesterday isn't really useful.

Today, hopefully after lunch, I will lock the PRMI with the new AC-coupled ASC servo, so that I can have the servo on, and the PRMI locked on the sideband, so that I have more light on the QPD. 

After that, it seems that the final thing we need to do before we vent is hold an arm near, but off resonance, lock the PRMI, and then swing the arm in and out of resonance a bit.

ELOG V3.1.3-